VT002

PI: Edwards
Title: VSOPT-TEST2

Session: June 1996 (session 2)
Frequency: 6 cm
Mode: VLBA
Obs time (UT): 168 EU 10:00(16/06)-15:00(16/06)
Stations: JbNtToShMh
Processor: Soc
EU Mk3 Tapes/St.: 1

Comment


To give feedback on this experiment, click here

Experiment Feedback

Torun (Roman Feiler at Sun Jun 16 10:53:24 EDT 1996) Minor failures
problems with tracking - 5 passes not recorded

Torun (Roman Feiler at Sun Jun 16 10:53:55 EDT 1996) Minor failures
problems with tracking - 5 passes not recorded

Metsahovi (Kaj Wiik at Sun Jun 16 16:24:13 EDT 1996) Minor failures
Our EFOS-9 Hmaser failed in friday morning and returned to operation in the afternoon. Offset was adjusted to 10 us in saturday evening. It seemed to work OK during the experiment. We received our new DSM from Penny&Giles just in time to install and test it. Except problems when operated without forced air cooling, it seemed to work satisfactorily. We missed some minutes from the start of the pass 9 (which started at 17:30 UT) because we were using the gap in the schedule to measure pointing and were late in restarting the schedule. We experienced a heavy rainfall during the first scan (3C84) but otherwise WX was reasonable.

Noto (Corrado Trigilio at Tue Jun 18 09:23:05 EDT 1996) Success
No problems reported. Tapes sent to Socorro on monday 17 june via CAI POST.

Jodrell (David Henstock at Fri Jun 21 09:03:13 EDT 1996) Minor failures
No problems; very successful. However a few minutes of data near start may be lost as the telescope was briefly off-source (the perils of having two field systems both able to control the telescope at the same time!) Tapes sent via Woodcroft Air Cargo (urgent delivery) on afternoon of Tuesday 18th.